# Tracking Brownian fluid particles in large eddy simulations

### v2

@@ 2.2 Representing the velocity in terms of fluid particles
$u(\xi,t)=\int_{\mathbb{R}^{d}}p_{u}(0,\eta;t,\xi)u_{0}(\eta)\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\left.g(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s$ (28)
@@
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta})\}}g(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s$ (29)
@@
We therefore have formulated the Navier-Stokes equations in terms of (29, 27, 19, 21). We shall use this formulation to implement large eddy simulations for viscous incompressible flows past a solid boundary.
3 Random LES
@@
$\chi(x)=\frac{1}{(2\pi\sigma^{2})^{\frac{d}{2}}}\exp\left(-\frac{|x|^{2}}{2\sigma^{2}}\right)\quad\textrm{ for }x\in\mathbb{R}^{d},$
@@ 3.1 Random LES for flows in $\mathbb{R}^{d}$
$\frac{\partial}{\partial t}\tilde{u}+\widetilde{(u\cdot\nabla)u}=\nu\Delta\tilde{u}-\nabla\tilde{p}+\tilde{F}$
@@
$\nabla\cdot\tilde{u}=0,$
@@
$\nabla\cdot\tau=(\tilde{u}\cdot\nabla)\tilde{u}-\widetilde{(u\cdot\nabla)u}$
@@
$\partial_{t}\tilde{u}+(\tilde{u}\cdot\nabla)\tilde{u}=\nu\Delta\tilde{u}-\nabla\tilde{p}+\tilde{F}-\nabla\cdot\tau.$
@@
$\tilde{u}(x,t)=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})g(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (31)
@@
$=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})g(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$
@@
Random LES method
@@
$\textrm{d}Y_{t}^{\eta}=U(Y_{t}^{\eta},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}^{\eta}=\eta$ (32)
@@
$U(x,t)=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-Y_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-Y_{t}^{\eta})G(Y_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (33)
@@
$G(x,t)=-\nabla P(x,t)+F(x,t)$ (34)
for every $x\in\mathbb{R}^{d}$ and $t\geq 0$, and $F=(F^{1},F^{2},F^{3})$ is an external force applying to the flow;
$\nabla P(x,t)=\int_{\mathbb{R}^{d}}K_{d}(x,y)\left.\left(\nabla\cdot F-\sum_{i,j=1}^{d}\frac{\partial U^{j}}{\partial y^{i}}\frac{\partial U^{i}}{\partial y^{j}}\right)\right|_{(y,t)}\textrm{d}y$ (35)
@@ 3.2 Random LES for flows past a wall
$\tilde{u}(x,t)=\int_{\mathbb{R}^{d}}\chi(x-y)u(y,t)\mathrm{d}y$
@@
$\tilde{u}^{i}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{+}(x,y)u^{i}(y,t)\mathrm{d}y,\quad\tilde{u}^{d}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{-}(x,y)u^{d}(y,t)\mathrm{d}y$ (36)
@@
$\tilde{u}^{i}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)u^{i}(y,t)\mathrm{d}y$
@@
$\;+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta,t-s<\zeta\circ\tau_{t}\right\}}\chi_{i}(x,X_{t}^{\eta})g^{i}(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$
@@
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta}\circ\tau_{t})\}}g^{i}(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,y)\textrm{d}\eta\textrm{d}s$ (38)
@@
$\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)u^{i}(y,t)$ $=\int_{\mathbb{R}_{+}^{d}}\left(\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)p_{u}(0,\eta;t,y)\textrm{d}y-\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)p_{u}(0,\bar{\eta};t,y)\textrm{d}y\right)u_{0}^{i}(\eta)\textrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta})\}}g^{i}(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,y)\textrm{d}y\textrm{d}\eta\textrm{d}s$ (39)
$=\int_{\mathbb{R}_{+}^{d}}\left(\mathbb{E}\left[\chi_{i}(x,X_{t}^{\eta})1_{\left\{t<\zeta\right\}}\right]-\mathbb{E}\left[\chi_{i}(x,X_{t}^{\bar{\eta}})1_{\left\{t<\zeta\right\}}\right]\right)u_{0}^{i}(\eta)\textrm{d}\eta$ (40)
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta\right\}}1_{\{t-s<\zeta\circ\tau_{t}\}}\chi_{i}(x,X_{t}^{\eta})g^{i}(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (41)
@@
Random LES method for flows past walls
@@
$\textrm{d}Y_{t}^{\eta}=U(Y_{t}^{\eta},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}^{\eta}=\eta$ (42)
@@
$\;+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta,t-s<\zeta\circ\tau_{t}\right\}}\chi_{i}(x,Y_{t}^{\eta})G^{i}(Y_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (43)
@@
$U(x,t)=\overline{U(\bar{x},t)}\quad\textrm{ for }x\in\mathbb{R}_{-}^{d},\quad U(x,t)=0\quad\textrm{ if }x_{d}=0,$ (44)
@@
$\;+\nu\int_{\mathbb{R}^{2}}\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,(y_{1},y_{2},0))\frac{\partial U^{2}}{\partial y_{3}}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$ (47)
@@
4 Numerical schemes and experiments
@@ 4.1 Numerical schemes and variables settings
Set mesh size $s>0$, time step $\delta>0$ and kinematic viscosity $\nu>0$. For $i_{1},i_{2},i_{3}\in\mathbb{Z}$, denote $y^{i_{1,}i_{2},i_{3}}=(i_{1},i_{2},i_{3})s$, $U^{i_{1},i_{2},i_{3}}=U(y^{i_{1,}i_{2},i_{3}},0)$. In numerical scheme, we drop the expectation and use one-copy of Brownian particles. We discrete the stochastic differential equation (42) following Euler scheme: for $t_{i}=i\delta,i=0,1,2,\cdots$.
$Y_{t_{k}}^{i_{1},i_{2},i_{3}}=Y_{t_{k-1}}^{i_{1},i_{2},i_{3}}+\delta U(Y_{t_{k-1}}^{i_{1},i_{2},i_{3}},t_{k-1})+\sqrt{2\nu}(B_{t_{k}}-B_{t_{k-1}}),\quad Y_{0}^{i_{1},i_{2},i_{3}}=y^{i_{1,}i_{2},i_{3}}$
where $Y_{t_{k}}^{i_{1},i_{2},i_{3}}$ denotes $Y_{t_{k}}^{x^{i_{1},i_{2},i_{3}}}$ for simplicity. Then integral representations in (43) and (47) can be discretized as follows:
$U(x,t_{k})$ $=\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}s^{3}\left[1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},i_{3}})-1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},-i_{3}})\right]U_{0}^{i_{1},i_{2},i_{3}}$
$+\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}\sum_{j=1}^{k}s^{3}\delta\prod_{l=j}^{k}1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},i_{3}})G(X_{t_{j-1}}^{i_{1},i_{2},i_{3}},t_{j-1})$ (48)
@@
$U(x,t)=\overline{U(\bar{x},t)}\quad\textrm{ for }x\in\mathbb{R}_{-}^{d},\quad U(x,t)=0\quad\textrm{ for }x_{3}=0,$ (49)
@@
$\nabla P(x,t_{k})$ $=\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}K_{3}^{+}(x,(i_{1},i_{2},i_{3}))\left(\nabla\cdot F((i_{1},i_{2},i_{3}),t_{k})-\sum_{i,j=1}^{3}\frac{\partial U^{j}}{\partial x^{i}}((i_{1},i_{2},i_{3}),t_{k})\frac{\partial U^{i}}{\partial x^{j}}((i_{1},i_{2},i_{3}),t_{k})\right)$
@@
$\;+\nu\sum_{i_{1},i_{2}\in\mathbb{R}}s^{2}\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,(i_{1},i_{2},0))\frac{\partial U^{2}}{\partial x_{3}}((i_{1},i_{2},0),t_{k})$ (50)
for $x=(x_{1},x_{2},x_{3})$ where $x_{3}>0$, the $\frac{\partial U^{j}}{\partial x^{i}}$ can be explicitly represented and calculated directly.
@@
In unbounded experiments, it is essential to introduce an initial vorticity into the fluid system to observe its subsequent evolution. This necessity arises because, in the absence of initial vorticity, an unbounded fluid system subjected solely to linear forces will not exhibit vortex dynamics. In contrast, in experiments involving wall-bounded systems, vorticity can naturally emerge due to boundary pressure effects, particularly near the walls. Furthermore, to facilitate the observation of vorticity formation, it is advisable to apply specific nonlinear forces concurrently. Therefore, in the subsequent experiments, we adopt the following setup: (1) for unbounded systems, an initial rotational velocity field is imposed along with linear forces; (2) for wall-bounded systems, an initial uniform velocity field (devoid of vorticity) is prescribed, accompanied by the application of nonlinear forces.
The experiments conducted without boundary conditions (4.2.1, 4.2.3, 4.3.1, 4.3.3) serve as a simplified "toy model" for preliminary exploration. Moving forward, our focus will shift to more realistic models that incorporate boundary conditions (4.2.2, 4.2.4, 4.3.2, 4.3.4). Through a comprehensive set of experiments, we aim to demonstrate the effectiveness of our approach.
4.2 Experiments for laminar flows
A laminar flow refers to a viscous flow characterized by a low Reynolds number. It is a smooth and orderly fluid motion in which the fluid moves in parallel layers without mixing between them. The mathematical description of laminar flow is relatively straightforward. Consequently, compared to turbulent flow, laminar flow requires less computational complexity and lower resource demands for numerical simulations.
4.2.1 2-dimensional unbounded laminar flows
In this experiment, we set $\nu=0.3$, $Re=1000$, length scale $L=3\pi$, so that $U_{0}=\frac{\nu}{L}Re=31.83$. The mesh size $s=\frac{3\pi}{50}\thicksim L\sqrt{\frac{1}{Re}}$ and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t_{1}=0.03,t_{2}=0.06,t_{3}=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0}\sin(2x_{2}),U_{0}\cos(2x_{1}))$, and set force $F=(10,-9.81)$. The velocity field and the vorticity field are shown in Figure 1.
Figure 1: Velocity fields (left) and vorticity (right) of unbounded laminar flows on $\mathbb{R}^{2}$
Since no external boundary conditions or nonlinear forces were applied in this experiment, an initial rotating vector field is necessary to observe the system’s evolution. As illustrated in Figure 1, the velocity field transitions from an initially rotating pattern to a more irregular state, eventually entering a chaotic regime. Correspondingly, the vorticity exhibits fluctuations that align with the expected behavior predicted by our experimental setup and hypothesis.
4.2.2 2-dimensional wall-bounded laminar flows

### v4

@@ 2.2 Representing the velocity in terms of fluid particles
$u(\xi,t)=\int_{\mathbb{R}^{d}}p_{u}(0,\eta;t,\xi)u_{0}(\eta)\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\left.g(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s,$ (32)
@@
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta})\}}g(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s,$ (33)
@@
See more proof processes and details in Appendix A. We therefore have formulated the Navier-Stokes equations in terms of (33, 31, 23, 25). We shall use this formulation to implement large-eddy simulation for viscous incompressible flows past a solid boundary.
3 Random LES method
@@
$\chi(x)=\frac{1}{(2\pi\sigma^{2})^{\frac{d}{2}}}\exp\left(-\frac{|x|^{2}}{2\sigma^{2}}\right)\quad\textrm{ for }x\in\mathbb{R}^{d},$ (34)
@@ 3.1 Random LES for flows in $\mathbb{R}^{d}$
$\frac{\partial}{\partial t}\tilde{u}+\widetilde{(u\cdot\nabla)u}=\nu\Delta\tilde{u}-\nabla\tilde{p}+\tilde{F}$, (36)
@@
$\nabla\cdot\tilde{u}=0.$ (37)
@@
$\nabla\cdot\tau=(\tilde{u}\cdot\nabla)\tilde{u}-\widetilde{(u\cdot\nabla)u}$. (38)
@@
$\partial_{t}\tilde{u}+(\tilde{u}\cdot\nabla)\tilde{u}=\nu\Delta\tilde{u}-\nabla\tilde{p}+\tilde{F}-\nabla\cdot\tau.$ (39)
@@
$\tilde{u}(x,t)=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})g(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s,$ (40)
@@
$=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})g(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s,$
@@
  • Numerical schemes for random LES method
@@
$\textrm{d}Y_{t}^{\eta}=U(Y_{t}^{\eta},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}^{\eta}=\eta,$ (41)
@@
$U(x,t)=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-Y_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-Y_{t}^{\eta})G(Y_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s,$ (42)
@@
$G(x,t)=-\nabla P(x,t)+F(x,t),$ (43)
for every $x\in\mathbb{R}^{d}$ and $t\geq 0$, $F=(F^{1},F^{2},F^{3})$ is an external force applying to the flow, and
$\nabla P(x,t)=\int_{\mathbb{R}^{d}}K_{d}(x,y)\left.\left(\nabla\cdot F-\sum_{i,j=1}^{d}\frac{\partial U^{j}}{\partial y^{i}}\frac{\partial U^{i}}{\partial y^{j}}\right)\right|_{(y,t)}\textrm{d}y,$ (44)
@@ 3.2 Random LES for flows past a wall
$\tilde{u}(x,t)=\int_{\mathbb{R}^{d}}\chi(x-y)u(y,t)\mathrm{d}y,$
@@
$\tilde{u}^{i}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{+}(x,y)u^{i}(y,t)\mathrm{d}y,\quad\tilde{u}^{d}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{-}(x,y)u^{d}(y,t)\mathrm{d}y,$ (45)
@@
$\tilde{u}^{i}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)u^{i}(y,t)\mathrm{d}y,$
@@
$\;+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta,t-s<\zeta\circ\tau_{t}\right\}}\chi_{i}(x,X_{t}^{\eta})g^{i}(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s,$ (47)
@@
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta}\circ\tau_{t})\}}g^{i}(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,y)\textrm{d}\eta\textrm{d}s,$
@@
$\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)u^{i}(y,t)$
$=\int_{\mathbb{R}_{+}^{d}}\left(\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)p_{u}(0,\eta;t,y)\textrm{d}y-\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)p_{u}(0,\bar{\eta};t,y)\textrm{d}y\right)u_{0}^{i}(\eta)\textrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta})\}}g^{i}(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,y)\textrm{d}y\textrm{d}\eta\textrm{d}s$
$=\int_{\mathbb{R}_{+}^{d}}\left(\mathbb{E}\left[\chi_{i}(x,X_{t}^{\eta})1_{\left\{t<\zeta\right\}}\right]-\mathbb{E}\left[\chi_{i}(x,X_{t}^{\bar{\eta}})1_{\left\{t<\zeta\right\}}\right]\right)u_{0}^{i}(\eta)\textrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta\right\}}1_{\{t-s<\zeta\circ\tau_{t}\}}\chi_{i}(x,X_{t}^{\eta})g^{i}(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s,$
@@
  • Numerical schemes for random LES method for flows past walls
@@
$\textrm{d}Y_{t}^{\eta}=U(Y_{t}^{\eta},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}^{\eta}=\eta,$ (48)
@@
$\;+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta,t-s<\zeta\circ\tau_{t}\right\}}\chi_{i}(x,Y_{t}^{\eta})G^{i}(Y_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s,$ (49)
@@
U(x,t)=U(x¯,t)¯ $\overline{U(\bar{x},t)}\quad\textrm{ for }x\in\mathbb{R}_{-}^{d},\quad U(x,t)=0\quad\textrm{ if }x_{d}=0$, (50)
@@
4 Numerical experiments and analysis
In this section, we present the numerical schemes in detail. Subsequently, we conduct comprehensive experiments to rigorously evaluate the effectiveness of the Random LES method. Finally, we provide an in-depth analysis of the proposed approach and perform systematic comparisons with existing methods, highlighting several distinct advantages of our technique.
@@ 4.1 Numerical schemes and variables settings
Set mesh size $s>0$, time step $\delta>0$ and kinematic viscosity $\nu>0$. For $i_{1},i_{2},i_{3}\in\mathbb{Z}$, denote $y^{i_{1,}i_{2},i_{3}}=(i_{1},i_{2},i_{3})s$, $U^{i_{1},i_{2},i_{3}}=U(y^{i_{1,}i_{2},i_{3}},0)$. In numerical scheme, we drop the expectation and use one-copy of Brownian particles. We can discrete the stochastic differential equation (48) following Euler scheme: for $t_{i}=i\delta,i=0,1,2,\cdots$.
$Y_{t_{k}}^{i_{1},i_{2},i_{3}}=Y_{t_{k-1}}^{i_{1},i_{2},i_{3}}+\delta U(Y_{t_{k-1}}^{i_{1},i_{2},i_{3}},t_{k-1})+\sqrt{2\nu}(B_{t_{k}}-B_{t_{k-1}}),\quad Y_{0}^{i_{1},i_{2},i_{3}}=y^{i_{1,}i_{2},i_{3}},$ (54)
where $Y_{t_{k}}^{i_{1},i_{2},i_{3}}$ denotes $Y_{t_{k}}^{x^{i_{1},i_{2},i_{3}}}$ for simplicity. However, the forward Euler method is not considered as an advanced and accurate enough method. To better calculate the stochastic differential equation, we use the Milstein method [20]. The Milstein method takes into account the derivative of the diffusion term and uses the Itô’s lemma expansion to correct the discretization error, the scheme is :
$Y_{t_{k}}^{i_{1},i_{2},i_{3}}$ $=Y_{t_{k-1}}^{i_{1},i_{2},i_{3}}+\delta U(Y_{t_{k-1}}^{i_{1},i_{2},i_{3}},t_{k-1})+\sqrt{2\nu}(B_{t_{k}}-B_{t_{k-1}})$
$+\frac{1}{2}U(Y_{t_{k-1}}^{i_{1},i_{2},i_{3}},t_{k-1})U^{\prime}(Y_{t_{k-1}}^{i_{1},i_{2},i_{3}},t_{k-1})[(B_{t_{k}}-B_{t_{k-1}})^{2}-\delta].$ (55)
The $U^{\prime}$ can be calculated explicitly. Then integral representations in (49) and (53) can be discretized as follows:
$U(x,t_{k})$ $=\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}s^{3}\left[1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},i_{3}})\right]U_{0}^{i_{1},i_{2},i_{3}}$
$-\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}s^{3}\left[1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},-i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},-i_{3}})\right]U_{0}^{i_{1},i_{2},i_{3}}$
$+\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}\sum_{j=1}^{k}s^{3}\delta\prod_{l=j}^{k}1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},i_{3}})G(X_{t_{j-1}}^{i_{1},i_{2},i_{3}},t_{j-1}),$ (56)
@@
U(x,t)=U(x¯,t)¯ $\overline{U(\bar{x},t)}\quad\textrm{ for }x\in\mathbb{R}_{-}^{d},\quad U(x,t)=0\quad\textrm{ for }x_{3}=0$, (57)
@@
$\nabla P(x,t_{k})$ $=\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}K_{3}^{+}(x,(i_{1},i_{2},i_{3}))\nabla\cdot F((i_{1},i_{2},i_{3}),t_{k})$
$-\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}K_{3}^{+}(x,(i_{1},i_{2},i_{3}))\sum_{i,j=1}^{3}\frac{\partial U^{j}}{\partial x^{i}}((i_{1},i_{2},i_{3}),t_{k})\frac{\partial U^{i}}{\partial x^{j}}((i_{1},i_{2},i_{3}),t_{k})$
@@
$\;+\nu\sum_{i_{1},i_{2}\in\mathbb{R}}s^{2}\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,(i_{1},i_{2},0))\frac{\partial U^{2}}{\partial x_{3}}((i_{1},i_{2},0),t_{k}),$ (58)
for $x=(x_{1},x_{2},x_{3})$ where $x_{3}>0$, the $\frac{\partial U^{j}}{\partial x^{i}}$ can be explicitly represented and calculated directly. Each variable is updated based on its value from the previous time step, the iterative computational procedure is as follows:
$U_{0},Y_{0}\ \rightarrow Y_{1}\xrightarrow{+G_{0},U_{0}}U_{1}\xrightarrow{+\frac{\partial U_{1}}{x},F}G_{1},$
$U_{1},Y_{1}\rightarrow Y_{2}\xrightarrow{+G_{1},U_{0}}U_{2}\xrightarrow{+\frac{\partial U_{2}}{x},F}G_{2},$
$\cdots\cdots$
$U_{t_{n-1}},Y_{n-1}\rightarrow Y_{T}\xrightarrow{+G_{t_{n-1}},U_{0}}U_{T}\xrightarrow{+\frac{\partial U_{T}}{x},F}G_{T}.$
@@
4.2 Experiments for random LES method
According to the numerical scheme above, we conducted four sets of experiments to simulate four distinct flow scenarios by setting different initial conditions and external forces.
  • Experiment 1: 2-dimensional laminar advective flow through a solid wall with external force
  • Experiment 2: 2-dimensional turbulent advective flow through a solid wall with external force
  • Experiment 3: 3-dimensional flow of an initially quiescent fluid driven by tensile forces.
  • Experiment 4: 3-dimensional turbulent advective flow through a solid wall with external force
4.2.1 Experiment 1: 2-dimensional advection through a wall (laminar flows)